The top part of Ramakrishnan, Inc,’s 2015 and 2014 balance sheets is listed below (in millions of dollars).
2015 2014 2015 2014
  Current assets:   Current liabilities:
      Cash and marketable securities $ 42 $ 33       Accrued wages and taxes $ 36 $ 35
      Accounts receivable 147 132       Accounts payable 91 80
      Inventory 210 191       Notes payable 72 64
           Total $ 399 $ 356            Total $ 199 $ 179

Calculate Ramakrishnan, Inc.’s current ratio for 2015 and 2014. (Round your answers to 2 decimal places.)

Current ratio
  2015 times
  2014 times

Calculate Ramakrishnan, Inc.’s quick ratio for 2015 and 2014. (Round your answers to 2 decimal places.)

Quick ratio
  2015 times
  2014 times

Calculate Ramakrishnan, Inc.’s cash ratio for 2015 and 2014. (Round your answers to 2 decimal places.)

Cash ratio
  2015 times
  2014 times
Answer #1

Calculate Ramakrishnan, Inc.’s current ratio for 2015 and 2014. (Round your answers to 2 decimal places.)

Current ratio
  2015 399/199 = 2.01 times
  2014 356/179 = 1.99 times

Calculate Ramakrishnan, Inc.’s quick ratio for 2015 and 2014. (Round your answers to 2 decimal places.)

Quick ratio
  2015 189/199 = 0.95 times
  2014 165/179 = 0.92 times

Calculate Ramakrishnan, Inc.’s cash ratio for 2015 and 2014. (Round your answers to 2 decimal places.)

Cash ratio
  2015 42/199 = 0.21 times
  2014 33/179 = 0.18 times